Circular No. 7953

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002ek W. M. Wood-Vasey, G. Aldering, P. Nugent, and R. Rudas, Lawrence

Berkeley National Laboratory, report the discovery of an apparent supernova (mag 19.2) on three unfiltered NEAT images taken with the 1.2-m Palomar Schmidt telescope (cf. SN 2002cx, IAUC 7902) on July 22.31, 22.33, and 22.37 UT. The new object is located at α = 20^{h}32^{m}08^{s}.50, δ = −15^{o}17′57′′.9 (equinox 2000.0), which is 4′′ west and 6′′ south of the center of the apparent host galaxy. SN 2002ek was confirmed on images taken on Aug. 6.27, 6.29, 6.32, 7.31, 7.33, and 7.36 at the same telescope, at mag 19.5 on both nights. No point source is visible at this location in previous images taken on during May 19–July 13 (limiting mag 21).

SUPERNOVA 2002el IN NGC 6986 B. Beutler and W. D. Li, University of California at Berkeley (UCB),

report the LOTOSS (cf. IAUC 7906) discovery of an apparent supernova on unfiltered KAIT images taken on Aug. 12.3 (mag ∼ 17.6) and 13.3 UT (mag ∼ 17.3). SN 2002el is located at α = 20^{h}56^{m}30^{s}.09, δ = −18^{o}33′34′′.3 (equinox 2000.0), which is 8′′.1 west and 24′′.8 north of the nucleus of NGC 6986. A KAIT image taken on Aug. 5.3 showed nothing at this position (limiting mag ∼ 18.5).

SUPERNOVAE 2002dr, 2002dx, 2002ea, 2002eb, 2002ee, 2002eh A. V. Filippenko, R. J. Foley, M. Papenkova, and B. Swift, UCB, com-

municate that inspection of CCD spectra (range 310–1000 nm), obtained on Aug. 9 UT with the Shane 3-m reflector at Lick Observatory, leads to the following spectral classifications of recent supernovae. SN 2002dr (IAUC 7927) is of type Ia, ∼ 0.5 year past maximum brightness, but the spectrum is noisy and heavily contaminated by light from the host galaxy. SN 2002dx (IAUC 7933) is also of type Ia, perhaps 2 months past maximum brightness; again, the spectrum is noisy and contaminated. SN 2002ea (IAUC 7936) is of type IIn, but with some low-velocity Hα absorption present. SN 2002eb (IAUC 7937) is of type Ia, close to maximum brightness. SN 2002eh (IAUC 7947) is also of type Ia, ∼ 1 month past maximum brightness. A spectrogram taken as above on Aug. 10 reveals that SN 2002ee (IAUC 7941) is a young type-II supernova, probably of the II-P variety. The continuum is quite blue, and well-developed P-Cyg profiles are visible, especially in the case of Hα.

2002 August 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green